OpenStudy (anonymous):

can someone solve this? (-7a^2+9u)-(-11a^2)

OpenStudy (anonymous):

u = 11/2

OpenStudy (anonymous):

I hope I helped you

OpenStudy (anonymous):

you have to combine like terms and solve it

OpenStudy (anonymous):

http://www.tiger-algebra.com/drill/-1=(9u-7/2)-(8u_9/3)/ this might help

OpenStudy (anonymous):

Use: \(\Large \color{red}{a-(-b)=a+b}\Rightarrow -(-11a^2)=11a^2\) \(\Large (-7a^2+9u)-(-11a^2)=-7a^2+9u+11a^2\\\Large =(11-7)a^2+9u=4a^2+9u\)

OpenStudy (anonymous):

thanks(:

OpenStudy (anonymous):

4a^2+9u